OpenStudy (anonymous):

peremeter is 40 the length is 4 ft longer than width1st equation is 2x+2y=?? I say 40, the what is the second equation x=y+??, then what is the length and what is the width

OpenStudy (anonymous):



OpenStudy (anonymous):

so you would have 2x+2y=40 then x=y+4 so what would be the number for the length and the width

OpenStudy (anonymous):

No...perimeter is all sides added together right? So w+w+w+4+w+4=40 Solve for w and youll have the width

OpenStudy (anonymous):

so you would do 40-8=32/4 right

OpenStudy (anonymous):

so the length would be 12 and the width would be 8 correct

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks so much for helping me, I like how people make it look like this and I figure it out in the long run thanks so much